OpenStudy (anonymous):

write an equation of the line that passes through the point (6 2) and is perpendicular to y= 3x + 3

OpenStudy (cj49):

slope is 3 y-y1=3(x-x1) y-2=3(x-6)

OpenStudy (anonymous):

thank you very much

OpenStudy (cj49):

yw c:...ull need to simplify it

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

\(\Huge\cal\color{blue}{?-?}\)

OpenStudy (triciaal):

just a minute the slope would not be 3

OpenStudy (triciaal):

parallel lines have the same slope but this line is perpendicular. for perpendicular lines the product of the slopes = -1

OpenStudy (cj49):

y=mx+c

OpenStudy (triciaal):

slope = -1/m= -1/3 points (6, 2) and (x, y) equation y-2 = -1/3(x - 6) point slope format y = -1/3 x + 4 slope intercept format

OpenStudy (triciaal):

@tabithachambers